# Diffusing States and Matching Scores: A New Framework for Imitation Learning

### v1

@@ Appendix A Error Analysis of Objective
In this section, we show the following two things: (1) why directly replacing the diffusion score function of $\pi$ with an estimator $g^{\pi}$ in $\ell(\pi)$ (Eq. 2) incurs large error; (2) why the error is small in our derived objective ^$\widehat{\ell}(\pi)$
@@
Appendix B Proof of Theorem 1
B.1 Supporting Lemmas
The following lemma is from Wang et al. (2023) and the fact that triangular discrimination is equivalent to the squared Hellinger distance up to a multiplicative constant: $2D_{H}^{2}\leq D_{\triangle}\leq 4D_{H}^{2}$(Topsoe, 2000).
@@ B.1 Supporting Lemmas
(Wang et al., 2023, Lemma 4.3) For two distributions $p,q\in[0,1]$ over some random variable, denote their respective means by $\overline{p}$ and $\overline{q}$. Then, it holds that
@@
The following lemma is adapted from some known results from diffusion processes and score matching. It shows that as long as the DS divergence (Definition 1) is small, the Hellinger distance is also small under some mild conditions. Similar results can be found in some theoretical papers on diffusion models such as Chen et al. (2022); Oko et al. (2023).
@@
First, by Item (a) and Girsanov’s Theorem (Karatzas & Shreve, 2014) (also see, e.g., Chen et al. (2022); Oko et al. (2023)), we have
@@
Next, by Item (b) and the convergence of OU process (e.g., Lemma 9 in Chen et al. (2023)), we have:
@@ B.3 Main proof
Appendix C Details of Experiments
@@
Table 2: Hyperparameters of our method and DAC. Shared columns indicate the same hyperparameters.

### v2

@@ Appendix B Error Analysis of Objective
In this section, we show the following two things: (1) why directly replacing the diffusion score function of $\pi$ with an estimator $g^{\pi}$ in $\ell(\pi)$ (Eq. 1) incurs large error; (2) why the error is small in our derived objective ^$\widehat{\ell}(\pi)$.
@@
Appendix C Proof of Theorem 1
C.1 Supporting Lemmas
The following lemma is from Wang et al. (2024c) and the fact that triangular discrimination is equivalent to the squared Hellinger distance up to a multiplicative constant: $2D_{H}^{2}\leq D_{\triangle}\leq 4D_{H}^{2}$(Topsoe, 2000).
@@ C.1 Supporting Lemmas
(Wang et al., 2024c, Lemma 4.3) For two distributions $p,q\in[0,1]$ over some random variable, denote their respective means by $\overline{p}$ and $\overline{q}$. Then, it holds that
@@
The following lemma is adapted from some known results from diffusion processes and score matching. It shows that as long as the DS divergence (Definition 1) is small, the Hellinger distance is also small under some mild conditions. Similar results can be found in some theoretical papers on diffusion models such as Chen et al. (2023b); Oko et al. (2023).
@@
First, by Item (a) and Girsanov’s Theorem (Karatzas & Shreve, 2014) (also see, e.g., Chen et al. (2023b); Oko et al. (2023)), we have
@@
Next, by Item (b) and the convergence of OU process (e.g., Lemma 9 in Chen et al. (2023a)), we have:
@@ C.3 Main proof
Appendix D Additional Experimental Results
In this section, we present additional experimental results: in Section D.1, we show the results of learning from state-action demonstrations; in Section D.2, we analyze the impact of the number of expert demonstrations on performance; in Section D.3, we study the expressiveness of diffusion models over other methods; in Section D.4, we include plots of unnormalized returns for reference; finally, in Section D.5, we provide detailed implementation details.
In addition, visual demonstrations of the experiments (such as time-lapse images) are provided in Figure 3 for better illustration.
D.1 Learning from State-action Demonstrations
D.2 On the Number of Expert Demonstrations
As we established in Section 5, our algorithm can have better sample complexity in terms of expert demonstrations. To empirically verify this, we conduct an experiment with varying numbers of expert demonstrations. We train SMILING and DAC with 1K, 2K, 5K, 10K, 25K, and 125K expert states on humanoid-walk and compare the results. The relationship between the number of expert demonstrations and the final performance (after 3M training steps) is shown in Figure 5 and the training curves of both algorithms are shown in Figure 6.
Figure 5: Relationship between the number of expert demonstrations and the final performance. The x-axis corresponds to the number of expert states. The y-axis is the final performance that is computed as the average of the last 100 training epochs. Each point is the average of five random seeds.
From Figure 5, we observe that SMILING dominates DAC across all expert demonstration sizes. In particular, SMILING is able to approach expert-level performance with 5K expert states, while DAC requires 25K to achieve similar performance. This suggests that SMILING is more efficient than DAC in utilizing expert demonstrations.
Figure 6: The training curves of SMILING (left) and DAC (right) with varying numbers of expert demonstrations. The x-axis corresponds to the number of environment steps (also the number of policy updates). The y-axis is the normalized return.
D.3 Expressiveness of Diffusion Models
In Appendix A, we argued that score matching is more expressive than discriminator-based methods, which may explain why our approach generally outperforms DAC in previous experiments. To further support this argument, we conducted an ablation study on the cheetah-run task where our method previously showed a slower convergence compared to DAC. Now we remove the activation functions in both the discriminator of DAC and the score function of our method, making them purely linear, and then re-run the experiments. The results are in Figure 7. Different from the previous experiments where DAC outperformed our method, in this case, DAC’s performance degrades notably by showing slower convergence and significantly higher variance. In comparison, our method remains more effective and outperforms DAC clearly.
We attribute the performance degradation of DAC to the limited expressiveness of the discriminator, which results in mode collapse—a well-known issue in discriminator-based methods, originally identified in GANs. In such methods, if the discriminator’s performance falls significantly behind the generator (e.g., due to insufficient training or a lack of discriminator expressiveness, which is the case here), the generator will exploit this gap and produce single-mode data that can easily fool the discriminator. In this case, the generator is falling into a local minima. As the discriminator gradually catches up and learns to identify this data, the generator will discover new data to fool the updated discriminator, falling into another local minimum. This perpetual cycle of chasing leads to continuous oscillations in the training process. We conjecture this is the exact reason for DAC’s instability in this experiments—With the discriminator reduced to a purely linear function, its expressiveness is significantly weakened, allowing the RL policy to exploit it easily, and thus a perpetual chasing emerges. However, our method can remain more effective even with linear score functions, which aligns with our hypothesis in Appendix A.
D.4 Plots with Unnormalized Returns
We provide the unnormalized-return versions (i.e., with the y-axis not normalized) of Figures 2 and 4 in Figures 8 and 9 for reference.
Figure 8: Same curves as in Figure 2 but with unnormalized return $V^{\pi}$ on the y-axis.
Figure 9: Same curves as in Figure 4 but with unnormalized return $V^{\pi}$ on the y-axis.
D.5 Implementation Details
We summarize the difficulties of the tasks evaluated in our experiments in Table 1.
@@ D.5 Implementation Details
Expert Training and Demonstration Collection. We use the aforementioned RL algorithms (SAC for DMC tasks and DreamerV3 for HumanoidBench tasks) to train the expert policies. We did not change the hyperparameters for these algorithms, making them the same as those provided in their original implementations in Yarats & Kostrikov (2020); Sferrazza et al. (2024). In other words, these algorithms were directly applied to their respective benchmarks without modification.
The RL algorithms were run until the performance plateaued, with the required number of training steps varying across tasks: ball-in-cup-catch (0.1M steps), cheetah-run (4.8M steps), humanoid-walk (3.4M steps), humanoid-sit (2.7M steps), humanoid-crawl (5.4M steps), and humanoid-pole (16M steps). The resulting expert policies achieve the following average total return: 972.85 for ball-in-cup-catch, 915.16 for cheetah-run, 875.61 for humanoid-walk, 794.12 for humanoid-sit, 914.31 for humanoid-crawl, and 879.48 for humanoid-pole.
Once the expert training is complete, we generate expert demonstrations by running the learned expert policy for five episodes for each task. Since each episode has a fixed horizon of 1K steps, the resulting dataset contains $5\times 1,000=5,000$ samples per task.
@@
Implementation of DiffAIL. We used the official implementation provided by Wang et al. (2024a) and made sure their neural networks have same size as ours. However, when applying their code to the tasks involving humanoids, we encountered numerical issues: the loss and weights of the neural networks diverged to infinity. We suspect this is due to numerical instability in their implementation, which becomes pronounced in the high-dimensional feature space of humanoid tasks.
Table 1: List of environments (“DMC” = DeepMind Control Suite, “HB” = HumanoidBench).
Task Difficulty Benchmark
Ball in Cup Catch Easy DMC
Cheetah Run Medium DMC
Humanoid Walk Hard DMC
Humanoid Sit Hard HB
Humanoid Crawl Hard HB
Humanoid Pole Hard HB
Table 2: Hyperparameters of our method and DAC. Shared columns indicate same hyperparameters.